A family needs a loan of 10 000 € for renovation. The loan period is 1 year and payments are made 4 times a year. The interest rate of the annuity loan is 4.2 % p.a. What is the remaining balance after the first payment?


1
Expert's answer
2022-02-02T15:40:54-0500

The family takes a loan of 10,000.

We have the following information.

"n=4\\i={0.042\over4}=0.0105" where, "n" is the loan period and "i" is the interest rate.

Let "P" be the quarterly installment. So "P" is given as,

"P={10,000\over{(1-1.0105^{-4})\over0.0105}}=2565.97"

"P=2565.97" is also the amount payable in the first installment.

The total payment after the 4 times of payment is, "2565.97\times 4=10263.87"

So the remaining balance after the first payment is, "10263.87-2565.97=7697.90"